Need help on this calculation sheet for my pre lab!
A determination of the molar solubility and the K_
Answer #1

5: Volume of HCl added = final buret reading - initial buret reading = 13.90 - 1.70 = 12.20 mL = 0.0122 L

6: Moles of HCl added = M x V(L) = 0.0480 mol/L x 0.0122 L = 5.856x10-4 mol

7: HCl + OH- ------ > H2O + Cl-

Hence moles of OH- in saturated solution = moles of HCl reacted = 5.856x10-4 mol

8: Total volume after neutralization, Vt = 25.00 mL + 12.20 mL = 37.20 mL = 0.0372 L

Hence equilibrium [OH-] = moles of OH- / Vt = 5.856x10-4 mol / 0.0372 L = 0.015742 mol/L

9: 1 mol of Ca(OH)2 contains 2 mol OH- and 1 mol Ca2+ ion. Hence

[Ca2+] = [OH-] / 2 =  0.015742 M / 2 = 0.007871 mol/L

10: Ca(OH)2 -----> Ca2+(aq) + 2OH-(aq)

-------------------------- S ------------- 2S

molar solubility, S = [Ca2+] = 0.007871 mol/L

11: Ksp = [Ca2+] x [OH-]2 = 0.007871 x (0.015742)2 = 1.95 x 10-5 (answer)